OpenStudy (anonymous):

You earn $245 in gross pay a week and 2/5 of your check is taken out for taxes and deductions. How much of your check is taken out for taxes and deductions?

OpenStudy (akashdeepdeb):

You earn $245. Now let us try to understand it: If you earn that much and 2/5ths of that is taken away, it means from every 5 in 245. 2 are taken away for taxes.... To understand it, I am going to show you one way. How many fives are there in 245? [245/5 = 49 fives are there] And thus FROM the 49 fives, 49 two's are taken away for taxes etc. 49 two's = 98$ That means 98$ from 245$ is taken away for taxes etc.... Did you understand this way? :D

OpenStudy (anonymous):

yes you're very helpful thank u

OpenStudy (anonymous):

Good explanation! Could also be by using fraction multiplication: \[\frac{ 245 }{ 1 }*\frac{ 2 }{ 5 } = \frac{490}{5} = 98\]

OpenStudy (anonymous):

... or rewrite the fraction 2/5 as a decimal number, since 2/5 = 0.4 = 40 %. 40 % of 245 is 245*0.4 = 98 :)

OpenStudy (akashdeepdeb):

:)